OpenStudy (anonymous):

Help me Find Lim sin(2-cos(x) - cos (2x) ) / x and x approaches zero

zepdrix (zepdrix):

Hey there :) Here is an helpful limit to recall:\[\large\rm \color{orangered}{\lim_{x\to0}\frac{\sin x}{x}=1}\]

zepdrix (zepdrix):

It might help us here, hmm thinking :d

zepdrix (zepdrix):

Hmm I can't find a normal approach to this one. Are you familiar with L'Hospital's Rule?

OpenStudy (anonymous):

No

zepdrix (zepdrix):

@Michele_Laino @Empty @AlexandervonHumboldt2 hmm any ideas people? :)

OpenStudy (michele_laino):

we can use this identity: \[\begin{gathered} \frac{{\sin \left( {2 - \cos x - \cos \left( {2x} \right)} \right)}}{x} = \hfill \\ \hfill \\ = \frac{{\sin \left( {2 - \cos x - \cos \left( {2x} \right)} \right)}}{{2 - \cos x - \cos \left( {2x} \right)}} \cdot \frac{{2 - \cos x - \cos \left( {2x} \right)}}{x} = \hfill \\ \hfill \\ = \frac{{\sin \left( {2 - \cos x - \cos \left( {2x} \right)} \right)}}{{2 - \cos x - \cos \left( {2x} \right)}} \cdot \frac{{1 - \cos x + 1 - \cos \left( {2x} \right)}}{x} = \hfill \\ \hfill \\ = \frac{{\sin \left( {2 - \cos x - \cos \left( {2x} \right)} \right)}}{{2 - \cos x - \cos \left( {2x} \right)}} \cdot \left\{ {\frac{{1 - \cos x}}{x} + \frac{{2\left( {1 - \cos \left( {2x} \right)} \right)}}{{2x}}} \right\} = \hfill \\ \hfill \\ = \frac{{\sin \left( {2 - \cos x - \cos \left( {2x} \right)} \right)}}{{2 - \cos x - \cos \left( {2x} \right)}} \cdot \left\{ {\frac{{2{{\left( {\sin \left( {x/2} \right)} \right)}^2}}}{x} + \frac{{4{{\left( {\sin x} \right)}^2}}}{x}} \right\} \hfill \\ \end{gathered} \]

OpenStudy (empty):

Looks like we're using the chain rule

OpenStudy (anonymous):

Thank you 😊